9- A simple random sample of dard deviation of $400. Construct a 9 36 college teachers has a meat of$1.000 (Monthly), with a sti- te 400. Construct a 99% confidence interval for the true mean salary < of college 15 pts) 10.- True/False Section (2pts cach) (a) For any data set we have (b) If Pi pthen 0 (c) If A and B are independent then P AnB PA)P(BA) (d) For any random sample to (e) For any random sample < P(x2 > d) (t) If P(A) 0, then P(A) st Q (h) For any two random samples (proportions) where n2 we have p - (i) If r 0 then t0 (i) For any two events we have PAnB) PAUB
Answer #1

Question 9

Mean monthly salary \bar x = $ 4,000

standard deviation = \sigma = $ 400

sample size n = 36

standard error of sample mean = se0 = 400/sqrt(36) = 400/6 = $ 66.67

99% confidence interval = \bar x +- Zcr se0 = 4000 +- 2.576 * 66.67 = ($ 3828.27, $ 4171.73)

Question 10

(a) False.

(b) False

(c) True

(d) True

(e) False

(f)True

(g) True

(h) False

(i) True

(j) True